OpenStudy (anonymous):

Let f(x) = 2x - 6. Solve f-1(x) when x = 2. -5 4 -2 7 Can anyone plz help me out? @texaschic101 @phi @Compassionate @radar @Loser66 @Preetha

OpenStudy (anonymous):

i got C

OpenStudy (anonymous):

do you know how to get the inverse function?

OpenStudy (phi):

you found f(x) which is different from f^-1(x)

OpenStudy (anonymous):

oh ok but idk how to find for the inverse.

OpenStudy (phi):

the way people are taught is to "swap x and y" in f(x) you start with y= 2x - 6 and swap x and y to get x= 2y - 6 if we "solve for y" we will find f^-1(x) we would add + 6 to both sides x+6= 2y then divide both sides by 2 (x+6)/2 = y now put in x=2 into that formula to find the answer.

OpenStudy (anonymous):

Thank you so much!!! @phi

OpenStudy (phi):

yw

OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6. Find f(g(x)). -40x + 24 -40x + 31 -40x + 64 -40x + 70 How would I solve for this one? @phi

OpenStudy (anonymous):

@kirbykirby can u plz help me out?

OpenStudy (anonymous):

@texaschic101 @phi @Compassionate @radar @Loser66 @Preetha